OpenStudy (anonymous):

Choose the correct simplification of the expression the quantity a to the 3rd power over b to the 7th power all raised to the 2nd power a to the 5th power over b to the 9th power a over b to the 14th power a to the 6th power over b to the 14th power a6b14

OpenStudy (johnweldon1993):

\[\large (\frac{ a^3 }{ b^7 })^2\] Like that?

OpenStudy (anonymous):

yes

OpenStudy (johnweldon1993):

So when the exponents are separated by the parenthesis...you multiply...and you apply it to each term...so \[\large (\frac{ a^3 }{ b^7 })^2 = \frac{ a^{3 \times 2} }{ b ^{7 \times 2} } = \frac{ a^{?} }{ b^{?} }\]

OpenStudy (anonymous):

a6 and b14

OpenStudy (johnweldon1993):

Right...just keep it as a fraction though...nothing changes in that sense...so we have: a to the 6th power over b to the 14th power

OpenStudy (anonymous):

cool thanks can you help one more time lol

OpenStudy (johnweldon1993):

Sure :)

OpenStudy (anonymous):

on this this. [7.04] Choose the correct simplification of the expression a to the 12th power times b to the 8th power all over a to the 2nd power times b to the 4th power. a10b4 a6b2 1 over a to the 10th power times b to the 4th power 1 over a to the 14th power times b to the 12th power

OpenStudy (johnweldon1993):

\[\large \frac{ a^{12} \times b^8 }{ a^2 \times b^4 }\] Like that I'm assuming...

OpenStudy (anonymous):

A96 and B8 but that's not right

OpenStudy (johnweldon1993):

When you are dividing exponents with the same base...you subtract the exponents and leave the base alone...so \[\large \frac{ a^{12}b^8 }{ a^2b^4 } = a^{12 - 2} \times b^{8 - 4} = a^{?}b^{?}\]

OpenStudy (anonymous):

A 10 B 4

OpenStudy (johnweldon1993):

Good! Which is answer choice...?

OpenStudy (anonymous):

it be A right

OpenStudy (johnweldon1993):

That would be correct!

OpenStudy (anonymous):

cool thanks you um can you still help me?????

OpenStudy (johnweldon1993):

Yeah sure if I can :)...and sorry was helping someone else

OpenStudy (anonymous):

its ok. [7.06] Choose the correct simplification of the expression −4x2(6x − 5x2 − 5). 20x4 + 24x3 + 20x2 −9x4 + 2x3 − 9x2 20x4 − 24x3 + 20x2 −20x4 + 24x3 − 20x2

OpenStudy (johnweldon1993):

So for this one...we take the term outside the parenthesis...(-4x^2) and multiply it by EVERY term INSIDE the parenthesis \[\large -4x^2 \times 6x = \space ?\] \[\large -4x^2 \times -5x^2 = \space ?\] \[\large -4x^2 \times -5 = \space ?\]

OpenStudy (anonymous):

20x4 − 24x3 + 20x2 it be this one right

OpenStudy (johnweldon1993):

That would be correct! Good work!

OpenStudy (anonymous):

thanks you can you help on this one too [7.07] Choose the correct product of (5x + 4)2. 25x2 + 40x + 16 25x2 − 16 25x2 + 16 25x2 − 40x + 16

OpenStudy (anonymous):

I think it will be this one but not sure lol = 25x^2 +40x +16

OpenStudy (johnweldon1993):

You are correct again! Pfft and you said you needed help :P

OpenStudy (anonymous):

will yea i am not good at math lol. [7.01] Choose the polynomial that is written in standard form. xy3 + 4x5y + 10x3 x4y2 + 4x2y + 8x −9x4y2 + 4x3y5 + 10x2 x6y3 + 4x4y8 + x2

OpenStudy (johnweldon1993):

What do you think it is first...?

OpenStudy (anonymous):

this one lol x^4y^2 + 4x^2y + 8x

OpenStudy (johnweldon1993):

Once again Correct! :)

OpenStudy (anonymous):

wow i am good for once in math lol [7.04] Choose the correct simplification of the expression (−9a3b2c10)0. 0 9a3b2c10 1 −9

OpenStudy (johnweldon1993):

Haha idk I think you were lying and are like some prodigy haha :) Your answer to this one is...?

OpenStudy (anonymous):

will it be 0 lol

OpenStudy (johnweldon1993):

HA finally I can help :P what is ANYTHING raised to the 0 power...?

OpenStudy (anonymous):

be 1

OpenStudy (johnweldon1993):

right....then your answer is?

OpenStudy (anonymous):

1 b/c If exponent is zero then the answer will always be one

OpenStudy (johnweldon1993):

That is correct in this case..because the 0 is outside the parenthesis so it gets applied to every term inside...so yes in this case...1

OpenStudy (anonymous):

cool help me on this b/c i don't know this one lol [7.08] Find the quotient of the quantity 21 times x to the 2nd power times y to the 6th power plus 6 times x times y to the 3rd power minus 30 times x times y all over 3 times x times y. 21x2y6 + 6xy3 − 10 7xy5 + 2y2 + 10 7xy5 + 2y2 − 10 7xy5 + 2y − 10

OpenStudy (johnweldon1993):

\[\large \frac{ 21x^2y^6 + 6xy^3 - 30xy}{ 3xy }\] Okay...this can just be broken up.....what is \[\large \frac{ 21x^2y^6 }{ 3xy } = \space ?\]

OpenStudy (anonymous):

7xy5+2y2-10

OpenStudy (johnweldon1993):

-_- yeah...you're a prodigy :P lol good work!

OpenStudy (anonymous):

thank you lol. [7.03] Identify the Exponent Law shown in the example below. x2x4 = x2+4 = x6 None of the choices below are correct. Power of a Product Power of a Power Product of Powers

OpenStudy (anonymous):

will it be this one Power of a product becuase your multiplying x^2 by x^4. But the powers add. Product of powers would be multiplying the powers. So x^2 by x^4 would be x^8

OpenStudy (johnweldon1993):

No this would be the product of powers rule...because you are starting with a product....but adding the powers....the power of a product would be something like (2x)^2 power of a power is something like (2x^2)^2 and the product of powers is this one

OpenStudy (anonymous):

oh ok. [7.08] Find the quotient of the quantity negative 12 times x to the 3rd power plus 21 times x to the 2nd power minus 6 times x all over negative 3 times x. −4x2 − 7x + 18 4x2 + 7x + 18 −12x3 + 21x2 + 2 4x2 − 7x + 2

OpenStudy (johnweldon1993):

mmhmm tell me the answer genius :P

OpenStudy (anonymous):

be this one −12x3 + 21x2 + 2 i am not sure lol

OpenStudy (johnweldon1993):

\[\large \frac{ -12x^3 + 21x^2 - 6x }{ -3x }\] what is \[\large \frac{ -12x^3 }{ -3x } = \space ?\]

OpenStudy (anonymous):

i don't know

OpenStudy (johnweldon1993):

What is -12 divided by -3?

OpenStudy (anonymous):

4

OpenStudy (johnweldon1993):

mmhmm...and what is x^3/x ? Remember when dividing exponents...you subtract themm...so x^3/x = x^(3 - 1) = x^?

OpenStudy (anonymous):

be this one 4x2 + 7x + 18 right

OpenStudy (johnweldon1993):

So close...but no the next step would be \[\large \frac{ 21x^2 }{ -3x }\] 21 / -3 = ?

OpenStudy (anonymous):

7

OpenStudy (johnweldon1993):

.....is it 7? remember you are dividing a positive by a negative....

OpenStudy (anonymous):

−4x2 − 7x + 18

OpenStudy (johnweldon1993):

When did the 4x^2 become a negative? that's not right either...

OpenStudy (johnweldon1993):

so far we have 4x² - 7x ....which one looks like that?

OpenStudy (anonymous):

4x2 − 7x + 2

OpenStudy (johnweldon1993):

There we go...you didn't seem too sure on that one...did you get it?

OpenStudy (anonymous):

yes i get it now. how about this one [7.04] Choose the correct simplification of the expression 4 times b all over a to the power of negative 10. a to the 10th power times b all over 4 4a10b 4 over a to the 10th power times b Already simplified

OpenStudy (anonymous):

will it be this one Already simplified

OpenStudy (johnweldon1993):

Not quite...we dont like negative exponents in the denominator... To make it positive...we simply make it go from the denominator...to the numerator...so we will have 4a^10b

OpenStudy (anonymous):

so it's B then??????

OpenStudy (johnweldon1993):

That would be correct :)

OpenStudy (anonymous):

cool :) [7.03] Choose the correct simplification of the expression (2x7y)2(y5)3. 4x7y10 4x9y10 4x14y16 4x14y17

OpenStudy (anonymous):

I think it will be 4x^14 y ^17?

OpenStudy (johnweldon1993):

Perfect!

OpenStudy (anonymous):

cool lol how about this one and i don't know how to do it Question 20 (Essay Worth 5 points) [7.02] Find the difference of (−5x2 − 4x3 − 3) − (9x3 + 3x − 8x2). Be sure to show all of your work in a neat and organized fashion. Use the ^ (shift 6) symbol to show an exponent. For example, x squared can be written as x^2.

OpenStudy (anonymous):

will be this Find the quotient of the quantity negative 5 times x to the 3rd power plus 20 times x to the 2nd power minus 25 times x all over negative 5 times x. (-5x^3+20x^2-25x)/(-5x)= x^2-4x+5

OpenStudy (anonymous):

Or this one First, convert the right side, since there's a negative sign in front of it. so it's -9x^3-3x+8x^2 Then put the whole expression into one long string: -5x^2-4x^3-3-9x^3-3x+8x^2 then start combing like terms: 3x^2 - 13x^3 - 3x -3 i am not sure lol

OpenStudy (johnweldon1993):

Okay....so for the "difference" one you posted... (-5x^2 - 4x^3 - 3) - (9x^3 + 3x - 8x^2) What we do first...is distribute the '-' sign through the 2nd parenthesis.... -5x^2 - 4x^3 - 3 -9x^3 - 3x + 8x^2 Now we just combine like terms... -4x^3 - 9x^3 - 5x^2 + 8x^2 - 3x - 3 -4x^3 - 9x^3 = -13x^3 -5x^2 + 8x^2 = 3x^2 so we have -13x^3 + 3x^2 - 3x - 3

OpenStudy (johnweldon1993):

The one you JUST posted is exactly what I just did

OpenStudy (anonymous):

lol that's funny lol :)

OpenStudy (johnweldon1993):

haha so that would be correct :)

OpenStudy (johnweldon1993):

All done now? :)